OpenStudy (anonymous):

HELP!? FAN AND MEDAL!! Margie wants to draw a pair of complementary angles. She drew one angel that had a measure of 35o. What should be the measure of the second angle? 1/ 35 2/ 55o 3/ 65o 4/ 145o Please say why it is the answer <3

OpenStudy (anonymous):

complementary angle has the sum of 90(degree) so the other angle must be 659degree)

OpenStudy (anonymous):

65 degree

OpenStudy (anonymous):



OpenStudy (anonymous):

35+A=90 A=90-35 A=65